OpenStudy (anonymous):

Find the derivative with respect to x of the given combination at the given value of x g(x+ f(x)), x = 3

OpenStudy (anonymous):

HOw is the answer -28?

OpenStudy (anonymous):

g(3+ 1) = g(4)

OpenStudy (anonymous):

Then I'm not sure what to do from there.

hartnn (hartnn):

first you need to differentiate...then put x=3, not before.

hartnn (hartnn):

whats derivative of g(x+f(x)) ??

OpenStudy (anonymous):

-4?

hartnn (hartnn):

using chain rule, \([g(x+f(x))]'=g'(x+f(x))[(x+f(x))]'\) got this ?

OpenStudy (anonymous):

Let me try to work it out using that formula... I actually don't remember seeing this before.

hartnn (hartnn):

chain rule is : \([g(f(x))]'= g'(f(x))[f'(x)]\)

OpenStudy (anonymous):

Still a little confused... [g(x+f(x))]′=g′(x+f(x))[(x+f(x))]′ = g'(4) * 4 ?

hartnn (hartnn):

lets go term by term you put x=3, f(3) =....? x+ f(x) = 3+f(3)=... ?

hartnn (hartnn):

i think u did that...and comes out to be 4

hartnn (hartnn):

g'(4) from the table =... ?

OpenStudy (anonymous):

g'(4) = -4

hartnn (hartnn):

yes, thats correct , now [x+f(x)]' = 1+ f'(x) = 1+ f'(3)=...?

OpenStudy (anonymous):

7

hartnn (hartnn):

how u got , [(x+f(x))]′ at x=3 as 4 ??

hartnn (hartnn):

you got how thats 7 ??

OpenStudy (anonymous):

now [x+f(x)]' = 1+ f'(x) = 1+ f'(3)=...? 1+ 6 = 7 ? f'(3) = 6

hartnn (hartnn):

yeah, i meant to ask, did you get how [x+f(x)]' = 1+ f'(x) ?

OpenStudy (anonymous):

Yeah, from the formula... The second part is [(x+f(x))]′ so we find that and multiply it by g′(x+f(x)) [g(x+f(x))]′=g′(x+f(x)) [(x+f(x))]′

OpenStudy (anonymous):

oh but why did x turn to 1 ?

hartnn (hartnn):

i didn't turn, i differentiated, \(\dfrac{d}{dx}(x)=1\)

hartnn (hartnn):

[x+f(x)]' = x'+f'(x) = 1+ f'(x)

hartnn (hartnn):

as i said, we first differentiate, then put x=.. any more doubts ?

OpenStudy (anonymous):

Ah okay. I get it now. You been very helpful! Thank you so much!

hartnn (hartnn):

welcome so much! :)